OpenStudy (anonymous):

PLEASE HELP! Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 A. (1, 1) B. (-1, -1) C. (1, -1) D. (-1, 1)

OpenStudy (anonymous):

The substitution method is where you try to get either \(x=\text{something to do with y}\) or \(y = \text{something to do with x}\), from one equation and substitute that information into the other equation. In this example, you may take equation 1\[2x+y = 3\], and rearrange it to become \[y=3-2x\]. Now in equation 2, replace \(y\) with \(3-2x\), and solve for x.

OpenStudy (anonymous):

Whats the answer? I still dont get it?

OpenStudy (anonymous):

Well if you take it a step further, where equation 2 normally looks like \[-3x+4y=1\], by performing this substitution, you replace \(y\) with \(3-2x\), so you get \[-3x + 4(3-2x)=1\]. Now you can use the various arithmetic and algebraic techniques in your toolbox to try to simplify this statement, and get down to the stage where you can say \[x=?\]